1.Monopolistic competition is a market structure in which

A.firms face perfectly elastic demand for their product.

B.firms face barriers to entry.

C.a large number of firms compete.

D.firms produce and sell an identical product.

E.the firms have no ability to influence the price of their product.

2.Which of the following is true of monopolistic competition in long−run ​equilibrium?

A.P​ > ATC and MR​ = MC

B.P​ = MR and P​ = MC

C.P​ > ATC and P​ > MR

D.P​ = ATC and P​ = MC

E.P​ = ATC and MR​ = MC

3. Suppose that Tommy​ Hilfiger's marginal cost of a jacket is a constant​ $100 and at one of the​ firm's shops, total fixed cost is​ $2,000 a day.  The​ profit-maximizing number of jackets sold in this shop is 20 a day. When the shops nearby start to advertise their​ jackets, this Tommy Hilfiger shop spends​ $2,000 a day advertising its​ jackets, and its​ profit-maximizing number of jackets sold jumps to 50 a day.

Calculate the​ shop's average total cost of a jacket sold before the advertising begins and after the advertising begins.

Before the advertising​ begins, the average total cost of a jacket in this shop is ​$____.

After the advertising​ begins, the average total cost of a jacket in this shop is ​$____.
Answer #1

1. C.a large number of firms compete.
(Demand curve is downward sloping, there is no barrier to entry, firms sell differentiated products, and have market power.)

2. E.P​ = ATC and MR​ = MC
(In long run, a monopolistically competitive firm earns zero economic profit, and MR = MC.)

3. Before the advertising,
TC = TFC + TVC = 2000 + (MC*Q) = 2000 + (100*20) = 2000+2000 = 4000
So, ATC = TC/Q = 4000/20
So, ATC = $200

After the advertising,
TC = TFC + TVC = 2000 + Advertising expenditiure + (MC*Q) = 2000 + 2000 + (100*50) = 4000+5000 = 9000
So, ATC = TC/Q = 9000/50
So, ATC = $180